how to get a positive seven answer using the integer negative 4 four times in an equation without using any other number?
Answer by Edwin McCravy(20054)


Absolute value of two additions and 1 division, in any order: 

|(-4)+(-4)+(-4)/(-4)| = 7

|(-4)+(-4)/(-4)+(-4)| = 7

|(-4)/(-4)+(-4)+(-4)| = 7

Edwin